OpenStudy (anonymous):

given that cotθ = 3, cos θ<0 find the exact value of cos θ/2

OpenStudy (anonymous):

I know to use the half angle for cosine but I don't know what to do with the cotangent.

OpenStudy (anonymous):

cotθ = 3 cos θ<0 ?

OpenStudy (anonymous):

cotθ =3 and cos θ<0

OpenStudy (campbell_st):

the angle is the the 3rd quadrant where tan is positive and cas is negative If cot = 3 then tan = 1/3 draw the diagram and find the hypotenuse using Pythagoras' theorem  \[\cos \theta = -\frac{3}{\sqrt{10}}\]